OpenStudy (anonymous):

find the slope of the line? -2 2 -1/2 1/2

OpenStudy (anonymous):

here is the slope.

OpenStudy (anonymous):

@Ogziii

OpenStudy (horsegirl27):

i can help!!

OpenStudy (anonymous):

question2, do the samething. find the slope of the line? -3 3 -1/3 1/3

OpenStudy (anonymous):

cool please help me

OpenStudy (horsegirl27):

Ok, do you know what slope is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol jk slope is a surface of which one end or side is at a higher level than another; a rising or falling surface.

OpenStudy (horsegirl27):

yup and remember, it's always rise over run. So u count up, then over

OpenStudy (anonymous):

we count slope this way , choose two points ( if the line is straigh ) x2 - x1 / y2 - y1 = 3 - 1 / 2 - ( -2 ) = 2 / 4 the slope is 1/2 hope that was helpul

OpenStudy (anonymous):

ok

OpenStudy (horsegirl27):

Thanks @Ogziii

OpenStudy (anonymous):

what about number 2

OpenStudy (anonymous):

is it 1/3

OpenStudy (horsegirl27):

You can do it the same way

OpenStudy (anonymous):

so but is it 1/3

OpenStudy (horsegirl27):

I think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the slope of a line that goes through (-3,-5) and (-3,-6) is. negative positive zero undefined

OpenStudy (anonymous):

@Ogziii

OpenStudy (anonymous):

the slope of a line that goes through (-3,-5) and (-3,-6) is. -3 -(-3) / -5 -(-6) = 0 / 1 = 0 the slope is zero so the function is constant

OpenStudy (anonymous):

for pic2.png whats the question ?

OpenStudy (anonymous):

for pic 2 the question is, find the slope of the line? -3 3 -1/3 1/3

OpenStudy (anonymous):

choose two points from the line , take their x's and y's , and then you use this x2 - x1 = y2 - y1 for example (2,1) and (1,4) 2 - 1 / 1 - 4 = -(1/3) it is simple !

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

will you were wrong i turned it in and all the questions are wrong except one the answer is actually 3